Daily requirements of 70 g of protein, 1 g calcium, 12 mg iron, and 3000 calories are needed for a balanced diet. The following foods are available for consumption with the cost and nutrients per 100 g as shown.

 

Protein

(g)

Calories

Calcium

(g)

Iron

Cost

GH¢

Brown Bread

12

246

0.1

3.2

0.5

Cheese

24.9

423

0.2

0.3

2

Butter

0.1

793

0.03

0

1

Baked Beans

6

93

0.05

2.3

0.25

Spinach

3

26

0.1

2

0.25

 

The objective is to find a balanced diet with minimum cost.

(a) Formulate a linear programming model for this problem.

(b) Use solver to find optimal solution and sensitivity report.     


1
Expert's answer
2021-03-15T19:31:16-0400

a) The objective is to find a balanced diet with minimum cost.

Let:

"x_1" - number of 100 g units of brown bread

"x_2" - number of 100 g units of cheese

"x_3" - number of 100 g units of butter

"x_4" - number of 100 g units of baked beans

"x_5" - number of 100 g units of baked beans

The linear programming problem is then:

Minimize: "0.5x_1 + 2x_2 + x_3 + 0.25x_4 + 0.25x_5"

Constraints:

"12x_1 + 24.9x_2 + 0.1x_3 + 6.0x_4 + 3.0x_5 \geq 70"

"246x_1 + 423x_2 + 793x_3 + 93x_4 + 26x_5 \geq 3000"

"0.1x_1 + 0.2x_2 + 0.03x_3 + 0.05x_4 + 0.1x_5 \geq 1"

"3.2x_1 + 0.3x_2 + 2.3x_4 + 2.0x_5 \geq 12"

"x_1,x_2,x_3,x_4,x_5\geq0"


b) Using online solver (https://cbom.atozmath.com), we get:


It was used Two-Phase Simplex method.

Phase 1:

after 1st step: maz "z_j-c_j=793.13"

after 2nd step: max "z_j-c_j=25.33"

after 3rd step: max "z_j-c_j=3.06"

after 4th step: max "z_j-c_j=0.075"

after 5th step: all "z_j-c_j\leq0"

optimal solution:

"min\ z=0,x_1=0.56,x_2=1.95,x_3=2.41,x_4=0,x_5=4.81"

Phase 1:

we eliminate the artificial variables and change the objective function for the original,

after 1st step: max "z_j-c_j=0.17"

after 2nd step: max "z_j-c_j=0.001"

after 3rd step: all "z_j-c_j\leq0"


Finally, optimal solution:

"z=\$5.64"

"x_1=9.77, x_3=0.75,x_2=x_4=x_5=0"


Conclusion:

To get minimal cost ("\$5.64" ) of diet, it's enough to use "977\ g" of brown bread and "75\ g" of butter. And we don't need any cheese, baked beans or spinach.